A bike costs $100.00 more dollars than a scooter. A scooter cost $60.00 more than a skateboard. The total cost of the 3 items is $310.00. How much is the skateboard? How much is the scooter? How much is the bike?
Answer by checkley79(3341)

B=s+100
S=s+60
B+S+s=310
s+100+s+60+s=310
3s+160=310
3s=310-160
3s=150
s=150/3
s=$50 ans. for the skateboard.
50+60=$110 ans. for the scooter
50+100=$150 ans for the bike.
Proof:
150+110+50=310
310=310